Afternoon tasks: ---------------- DA and instrument checkout completed with no problems. Jon & Warren worked some more this afternoon on the secondary galil motor connections. They replaced pins and replaced spliced wires. Steph ran the secondary back & forth several times, but could not reproduce the "not stopping on a full step error". Let's hope this is fixed. Robert & Steph verified that the same fiducial tables were used for both pointing models taken this dark run. The difference between the two models that was noted last night is not due to the fiducial tables. Pete & Craig called Russell to see if there were any TCC related problems that could have caused the difference. PT: --- Biases and dome flats taken.